The diameters of washers are normally distributed with mean 50mm and variance



4mm. A washer is considered non-defective if its diameter lies between 46mm and



53mm. If one washer is randomly selected, what is the probability that:



4.2.1 its diameter less than 45mm? (6)



4.2.2 itis defective?

1
Expert's answer
2022-06-08T16:26:05-0400

Let "X=" diameters of washer: "X\sim N(\mu, \sigma^2)."

Given "\mu=50mm, \sigma^2=4mm^2"

4.2.1


"P(X<45)=P(Z<\dfrac{45-50}{\sqrt{4}})"

"=P(Z<-2.5)\approx0.006210"

4.2.2


"P(46<X<53)=P(Z<\dfrac{53-50}{\sqrt{4}})"

"-P(Z\le\dfrac{46-50}{\sqrt{4}})=P(Z<1.5)-P(Z\le-2)"

"\approx0.9331928-0.0227501"

"\approx0.910443"

"P(defective)=1-0.910443=0.089557"